OpenStudy (anonymous):

Find the Taylor polynomial of degree n approximating the given function for x near 0. sqrt(1+x) How do I solve these sorts of problems

ganeshie8 (ganeshie8):

taylor series at x = 0 gives u maclaurin series

ganeshie8 (ganeshie8):

\(\large f(x) = f(0) + xf'(0) + \frac{x^2}{2!}f''(0) + \frac{x^3}{3!}f'''(0) + ....\)

ganeshie8 (ganeshie8):

start by finding first "n" derivatives of given function and plug them above

ganeshie8 (ganeshie8):

\(\large f(x) = \sqrt{1+x}\) \(\large f(0) = ?\) \(\large f'(0) = ?\) \(\large f''(0) = ?\) \(\large f'''(0) = ?\)

OpenStudy (anonymous):

ohh okay, so 1/2 -1/4 3/8 and so forth. thanks! :)

ganeshie8 (ganeshie8):

yes, divide them by factorial terms also that gives coefficients of polynomial